# Stacky Batyrev–Manin conjecture and modular curves

### v1

@@ 3. Modular curves
We now treat the general case. The stack $\mathscr{X}_{0}(N)$ has the untwisted sector, that we denote by 0 and which has age 0. Let us choose $x\in\mathscr{X}_{0}(N)(\overline{F})$. Suppose that under $J_{N}:\mathscr{X}_{0}(N)\to\mathscr{X}_{0}(1)$, the point $x$ does not map to the points $j=0$ or $j=1728$ of $\mathscr{X}_{0}(1)$. Let $U$ be the complement of the images of $j=0$ and $j=1728$ in $\mathbb{P}^{1}$. The automorphism group of $x$ is isomorphic to $\mu_{2}$. We deduce that $x$ maps to a non-stacky point of $\mathscr{X}_{0}(N)^{\operatorname{rig}}$. Let $s_{2}:\widehat{\mu}\to\mu_{2}$ be the canonical surjection defined over $\operatorname{Spec}(\mathbb{Z})$. The points $(x,\widehat{\mu}\xrightarrow{x\mapsto 1}\mu_{2})$ and $(x,\widehat{\mu}\xrightarrow{(s_{2})_{\overline{F}}}\mu_{2})$ of $\mathcal{J}_{0}\mathscr{X}_{0}(N)(\overline{F})$ map to points of $\mathcal{J}_{0}\mathscr{X}_{0}(1)(\overline{F})$ which lie in distinct connected components. It follows that they lie in distinct connected components of $\mathcal{J}_{0}\mathscr{X}_{0}(N)$. The point $(x,\widehat{\mu}\xrightarrow{x\mapsto 1}\mu_{2})$ lies in the sector 0. We show that $(x_{1},\widehat{\mu}\xrightarrow{(s_{2})_{\overline{F}}}\mu_{2})$ and $(x_{2},\widehat{\mu}\xrightarrow{(s_{2})_{\overline{F}}}\mu_{2})$, with $x_{1}$ and $x_{2}$ not mapping to $j=0,1728$, lie in the same component. Indeed, let $\overline{x_{1}}$ and $\overline{x_{2}}$ be the images in $U_{\overline{F}}$ of $x_{1}$ and $x_{2}$. We have that $\mathscr{X}_{0}(N)|_{U}\cong U\times B\mu_{2}$ because $\operatorname{Pic}(U)=\{0\}$. Consider $(U\xrightarrow{(\operatorname{Id}_{U},1)}U\times B\mu_{2}\to\mathscr{X}_{0}(N),((\widehat{\mu})_{U}\xrightarrow{(s_{2})_{U}}(\mu_{2})_{U})$. Restricting it to $\overline{x_{i}}$ gives $(x_{i},\widehat{\mu}\xrightarrow{(s_{2})_{\overline{F}}}\mu_{2})$ for $i=1,2$, which means that they are in the same connected component. We call this sector $1/2$. Now it maps to the sector $1/2$ of $\mathscr{X}_{0}(1)$, and Corrolary 2.13 implies that its age is 0.
@@
Consider now a geometric point $x$ above $j=0$ with automorphism group scheme $\mu_{2}$. Let $S$ be the closed substack of $\mathscr{X}_{0}(N)$ with automorphism group scheme larger than $\mu_{2}$. Note that $S$ is finite, i. e. $S\langle\overline{F}\rangle$ is finite. Let $y$ be another finite closed substack of $\mathscr{X}_{0}(N)$ which is geometrically a singleton, distinct from $x$ and distinct from $S$. Set $S^{\prime}:=S\cup\{y\}$. One has that $V:=(\mathscr{X}_{0}(N)-S^{\prime})^{\operatorname{rig}}$ is an open subscheme of $\mathbb{P}^{1}$ and not equal to $\mathbb{P}^{1}$. Hence $\operatorname{Pic}(V)=\{0\}$ and $\mathscr{X}_{0}(N)|_{V}\cong V\times\mathbb{P}^{1}$. Now as above, we have that $(x,(\widehat{\mu})_{\overline{F}}\xrightarrow{s_{\overline{F}}}(\mu)_{\overline{F}})$ is in the sector $1/2$. The same arguments apply if $x$ is a point with automorphism group $\mu_{2}$ lying above $j=1728$.

### v2

@@ 3. Modular curves
We now treat the general case. The stack $\mathscr{X}_{0}(N)$ has the untwisted sector, denoted by 0, whose age is 0. Choose $x\in\mathscr{X}_{0}(N)(\overline{F})$. Suppose that $j(J_{N}(x))\neq 0,1728$. Let $U$ be the complement of $j=0$ and $j=1728$ in $\mathbb{P}^{1}\cong X_{0}(1)$. Then $\operatorname{Aut}(x)$ is $\mu_{2}$. In this case, $x$ maps to a non-stacky point of $\mathscr{X}_{0}(N)^{\operatorname{rig}}$. Let $s_{2}:\widehat{\mu}\to\mu_{2}$ be the canonical surjection defined over $\operatorname{Spec}(\mathbb{Z})$. The points $(x,\widehat{\mu}\xrightarrow{x\mapsto 1}\mu_{2})$ and $(x,\widehat{\mu}\xrightarrow{(s_{2})_{\overline{F}}}\mu_{2})$ of $\mathcal{J}_{0}\mathscr{X}_{0}(N)(\overline{F})$ map to points of $\mathcal{J}_{0}\mathscr{X}_{0}(1)(\overline{F})$ which lie in distinct connected components, so they lie in distinct connected components of $\mathcal{J}_{0}\mathscr{X}_{0}(N)$. The point $(x,\widehat{\mu}\xrightarrow{x\mapsto 1}\mu_{2})$ lies in the sector 0. We show that $(x_{1},\widehat{\mu}\xrightarrow{(s_{2})_{\overline{F}}}\mu_{2})$ and $(x_{2},\widehat{\mu}\xrightarrow{(s_{2})_{\overline{F}}}\mu_{2})$, with $x_{1}$ and $x_{2}$ not mapping to $j=0,1728$, lie in the same component. Indeed, let $\overline{x_{1}}$ and $\overline{x_{2}}$ be the images in $U_{\overline{F}}$ of $x_{1}$ and $x_{2}$. We have that $\mathscr{X}_{0}(N)|_{U}\cong U\times B\mu_{2}$ because $\operatorname{Pic}(U)=\{0\}$. Consider $(U\xrightarrow{(\operatorname{Id}_{U},1)}U\times B\mu_{2}\to\mathscr{X}_{0}(N),((\widehat{\mu})_{U}\xrightarrow{(s_{2})_{U}}(\mu_{2})_{U})$. Restricting it to $\overline{x_{i}}$ gives $(x_{i},\widehat{\mu}\xrightarrow{(s_{2})_{\overline{F}}}\mu_{2})$ for $i=1,2$, which means that they are in the same connected component. We call this sector $1/2$. Now it maps to the sector $1/2$ of $\mathscr{X}_{0}(1)$, and Corrolary 2.13 implies that its age is 0.
@@
Consider now a geometric point $x$ above $j=0$ with automorphism group scheme $\mu_{2}$. Let $S$ be the closed substack of $\mathscr{X}_{0}(N)$ with automorphism group scheme larger than $\mu_{2}$. Note that $S$ is finite, i. e. $S\langle\overline{F}\rangle$ is finite. Let $y$ be another finite closed substack of $\mathscr{X}_{0}(N)$ which is geometrically a singleton, disjoint from $\{x\}\cup S$. Set $S^{\prime}:=S\cup\{y\}$. One has that $V:=(\mathscr{X}_{0}(N)-S^{\prime})^{\operatorname{rig}}$ is an open subscheme of $\mathbb{P}^{1}$ and not equal to $\mathbb{P}^{1}$. Hence $\operatorname{Pic}(V)=\{0\}$ and $\mathscr{X}_{0}(N)|_{V}\cong V\times\mathbb{P}^{1}$. Now as above, we have that $(x,(\widehat{\mu})_{\overline{F}}\xrightarrow{s_{\overline{F}}}(\mu)_{\overline{F}})$ is in the sector $1/2$. The same arguments apply if $x$ is a point with automorphism group $\mu_{2}$ lying above $j=1728$.